OpenStudy (anonymous):

What is the square root of this number? 48

OpenStudy (anonymous):

6.92820323028

OpenStudy (anonymous):

6.928

OpenStudy (anonymous):

thats no one of my answers

OpenStudy (anonymous):

6.928

OpenStudy (anonymous):

\[\sqrt{48}= \sqrt{16 \times 3} = 4\sqrt3\]

OpenStudy (anonymous):

maybe you need to round it to 7?

OpenStudy (anonymous):

Is there a 7 in the answer options? Maybe they want you to round it to 7.

OpenStudy (anonymous):

what are your answers?

OpenStudy (anonymous):

or maybe \[4\sqrt{3}\]

OpenStudy (anonymous):

4 times the square root of 3 might be it

OpenStudy (anonymous):

@twilightrain don't be rude.